Today, your paint line must produce 6250 gallons of beige paint. Your four step manufacturing process has the following scrap rates for the corresponding step:

Step 1: 0.4%

Step 2: 1.4%

Step 3: 2.3%

Step 4: 1.3%

How many gallons of paint must you start with to have 6250 gallons of point that can be sold?

(Keep one decimal point in your answer)
Answer #1

Answer: we need to input 6599.8 gallons

Explanation:

44 B C D E F G
45 Scrap rates 0.40% 1.40% 2.30% 1.30%
46 Input step-1 step-2 step-3 step-4 Output
47 6599.8 6599.8 6573.4 6481.4 6332.3 6250

Excel formula

44 B C D E F G
45 Scrap rates 0.004 0.014 0.023 0.013
46 Input step-1 step-2 step-3 step-4 Output
47 =C47 =D47/(1-C45) =E47/(1-D45) =F47/(1-E45) =G47/(1-F45) 6250